Write an equation of the line that passes through (-3,4) and is parallel to y=-3x-1
Answer by sachi(548)

the equation of the line that passes through (-3,4)
and is parallel to y=-3x-1 i.e slope=-3
is [y-4]/[x-(-3)]=-3
or y-4=-3(x+3)=-3x-9
or 3x+y+5=0